OpenStudy (anonymous):

ok find a1 a2 a3 a4 when you have a subscript n = 2(n+3)!

OpenStudy (anonymous):

\[a_n=2(n+3)!\]

OpenStudy (anonymous):

replace n by 1, 2, 3 just like a function. for example \[a_1=2(1+3)!\] \[a_1=2\times 4!\] \[a_1=2\times 4\times 3\times 2=48\]